# Robust First and Second-Order Differentiation for Regularized Optimal Transport

### v1

@@ 3.2 Analytical computation of the gradients
where $\mathbf{e}_{k}$ is the $k$-th standard column basis of $\mathbb{R}^{M}$ and the matrix $B_{k}\in\mathbb{R}^{d\times N}$ is $(B_{k})_{sj}=\frac{\partial C_{kj}}{\partial(\boldsymbol{y}_{k})_{s}}\Pi^{*}_{kj}$. Instead of solving the above linear system to evaluate (16a), we observe from the marginal constraints that
$\left[\begin{array}[]{ccccccccc}\frac{1}{M}\mathbbm{1}^{\top}_{M}&\frac{1}{N}\mathbbm{1}^{\top}_{N}\end{array}\right]H(\Pi^{*})=\left(\frac{1}{M}+\frac{1}{N}\right)\left[\begin{array}[]{ccccccccc}\boldsymbol{\mu}^{\top}&\boldsymbol{\nu}^{\top}\end{array}\right].$ (29)
Then, multiplying both sides of (26) by $\left[\begin{array}[]{ccccccccc}\frac{1}{M}\mathbbm{1}^{\top}_{M}&\frac{1}{N}\mathbbm{1}^{\top}_{N}\end{array}\right]$, we obtain
@@
$\left[\begin{array}[]{ccccccccc}\frac{1}{M}\mathbbm{1}^{\top}_{M}&\frac{1}{N}\mathbbm{1}^{\top}_{N}\end{array}\right]H(\Pi^{*})\left[\begin{array}[]{ccccccccc}\frac{d\boldsymbol{f}^{*}}{d\boldsymbol{y}_{k}}\\ \frac{d\boldsymbol{g}^{*}}{d\boldsymbol{y}_{k}}\end{array}\right]$ $\left[\begin{array}[]{ccccccccc}\frac{d\boldsymbol{f}^{*}}{d\boldsymbol{Y}}\\ \frac{d\boldsymbol{g}^{*}}{d\boldsymbol{Y}}\end{array}\right]$ $=\left[\begin{array}[]{ccccccccc}\frac{1}{M}\mathbbm{1}^{\top}_{M}&\frac{1}{N}\mathbbm{1}^{\top}_{N}\end{array}\right]\left[\begin{array}[]{ccccccccc}\mathbf{e}_{k}(B_{k}\mathbbm{1}_{N})^{\top}\\ B_{k}^{\top}\end{array}\right].$
Hence, the derivative of $\mathrm{OT}_{\epsilon}(C,\boldsymbol{\mu},\boldsymbol{\nu})$ with respect to source data $\boldsymbol{y}_{k}$ is
@@
$\frac{d\mathrm{OT}_{\epsilon}(C,\boldsymbol{\mu},\boldsymbol{\nu})}{d\boldsymbol{y}_{k}}=\left[\begin{array}[]{ccccccccc}\boldsymbol{\mu}^{\top}&\boldsymbol{\nu}^{\top}\end{array}\right]\left[\begin{array}[]{ccccccccc}\frac{d\boldsymbol{f}^{*}}{d\boldsymbol{y}_{k}}\\ \frac{d\boldsymbol{g}^{*}}{d\boldsymbol{y}_{k}}\end{array}\right]$
@@
$=\frac{1}{\frac{1}{M}+\frac{1}{N}}\left[\begin{array}[]{ccccccccc}\frac{1}{M}\mathbbm{1}^{\top}_{M}&\frac{1}{N}\mathbbm{1}^{\top}_{N}\end{array}\right]\left[\begin{array}[]{ccccccccc}\mathbf{e}_{k}(B_{k}\mathbbm{1}_{N})^{\top}\\ B_{k}^{\top}\end{array}\right]B_{k}\mathbbm{1}_{N}=\sum_{j=1}^{N}\frac{\partial C_{kj}}{\partial\boldsymbol{y}_{k}}\Pi^{*}_{kj}.$
@@
If we define the third-order tensor $\mathcal{B}$ by stacking the matrices $\{B_{k}\}_{k=1}^{M}$, that is the $k$-th component $\mathcal{B}_{k\cdot\cdot}=B_{k}$, then the linear system (26) can be further vectorized as
@@
$H(\Pi^{*})\left[\begin{array}[]{ccccccccc}\frac{d\boldsymbol{f}^{*}}{d\boldsymbol{Y}}\\ \frac{d\boldsymbol{g}^{*}}{d\boldsymbol{Y}}\end{array}\right]=\mathcal{R},\qquad\text{with}\quad\mathcal{R}:=\left[\begin{array}[]{ccccccccc}\text{diag}(\mathcal{B}\cdot\mathbbm{1}_{N})\\ \mathcal{B}^{\top}\end{array}\right]\in\mathbb{R}^{(M+N)\times M\times d}$ (34)
@@
where $\text{diag}(\mathcal{B}\cdot\mathbbm{1}_{N})$ is a third-order tensor, with $\text{diag}(\mathcal{B}\cdot\mathbbm{1}_{N})_{kks}=(\mathcal{B}\cdot\mathbbm{1}_{N})_{ks}$ and zeros for the other entries, and $\mathcal{B}^{\top}$ is the transpose of permutation of indices defined by $\left(\mathcal{B}^{\top}\right)_{ijk}=\mathcal{B}_{kij}$. Hence, by plugging (34) back to (16b), the derivative of $\text{OT}_{\epsilon}(C,\boldsymbol{\mu},\boldsymbol{\nu})$ with respect to the source data $\boldsymbol{Y}$ in the tensor form is equal to $\frac{d\mathrm{OT}_{\epsilon}(C,\boldsymbol{\mu},\boldsymbol{\nu})}{d\boldsymbol{Y}}=\mathcal{B}\cdot\mathbbm{1}_{N}$. ∎
@@
This theorem not only implies that automatic differentiation is unnecessary for computing the gradient of EOT distance but also that solving the large linear system (34) for $\frac{d\boldsymbol{f}^{*}}{d\boldsymbol{y}_{k}}$ and $\frac{d\boldsymbol{g}^{*}}{d\boldsymbol{y}_{k}}$ is not needed. The analytical expression of the gradient follows directly from the cost matrix $C_{\theta}$ and optimal coupling matrix $\Pi^{*}$ from the Sinkhorn algorithm, which can significantly speed up the computation. This theorem provides a different approach to derive the same analytical result other than the Danskin’s theorem. We emphasize that the gradient formula (18) is generic and applies to any choice of cost matrix $C$ without requiring $\boldsymbol{\mu}$ and $\boldsymbol{\nu}$ to be uniform. In particular, if the Sinkhorn iteration stops early and the coupling matrix $\widehat{\Pi}$ is suboptimal, the formula is still exact for $\frac{d\mathrm{OT}_{\epsilon}(C,\tilde{\boldsymbol{\mu}},\tilde{\boldsymbol{\nu}})}{d\boldsymbol{Y}}$ given suboptimal coupling matrix $\widehat{\Pi}$ and associated marginals $\tilde{\boldsymbol{\mu}}$ and $\tilde{\boldsymbol{\nu}}$, where $\sum_{j}\widehat{\Pi}_{ij}=\widetilde{\mu}_{i}$ and $\sum_{i}\widehat{\Pi}_{ij}=\widetilde{\nu}_{j}$.
@@
Gradient of $\widetilde{\text{OT}}_{\epsilon}(C,\boldsymbol{\mu},\boldsymbol{\nu})$. The above techniques also applies for computing the gradient of the Sinkhorn distance $\widetilde{\text{OT}}_{\epsilon}(C,\boldsymbol{\mu},\boldsymbol{\nu})$. Although it does not yield an analytical expression, it helps significantly reduce the computational cost by avoiding directly solving the linear system (34), which is costly since the right-hand-side is a large third-order tensor $\mathcal{R}$. Specifically, we have the following novel result.
@@
The gradient of Sinkhorn distance in (12) with respect to the source data $\boldsymbol{Y}$ is
$\frac{d\widetilde{\mathrm{OT}}_{\epsilon}(C,\boldsymbol{\mu},\boldsymbol{\nu})}{d\boldsymbol{Y}}=\left(\mathcal{B}-\frac{\mathcal{A}}{\epsilon}\right)\cdot\mathbbm{1}_{N}+\frac{1}{\epsilon}\boldsymbol{r}^{\top}\mathcal{R},$ (35)
where $\mathcal{B}$ and $\mathcal{R}$ are defined in (34), the third-order tensor $\mathcal{A}$ is $\mathcal{A}_{ksj}=\frac{\partial C_{kj}}{\partial(\boldsymbol{y}_{k})_{s}}C_{kj}\Pi^{*}_{kj}$. The vector $\boldsymbol{r}$ is the solution of the linear system, $H(\Pi^{*})\boldsymbol{r}=\left[\begin{array}[]{ccccccccc}\boldsymbol{a}\\ \boldsymbol{b}\end{array}\right]$, where vectors $\boldsymbol{a}$ and $\boldsymbol{b}$ have entries $a_{i}=\sum_{j=1}^{M}C_{ij}\Pi^{*}_{ij}$ and $b_{j}=\sum_{i=1}^{N}C_{ij}\Pi^{*}_{ij}$.
@@
From (12) and (19), the derivative of Sinkhorn distance with respect to the data ${y}_{k}$ is
$\frac{d\widetilde{\mathrm{OT}}_{\epsilon}(C,\boldsymbol{\mu},\boldsymbol{\nu})}{d\boldsymbol{y}_{k}}$ $=\sum_{j=1}^{N}\left(1-\frac{C_{kj}}{\epsilon}\right)\Pi^{*}_{kj}\frac{\partial C_{kj}}{\partial\boldsymbol{y}_{k}}+\frac{1}{\epsilon}\left(\sum_{i=1}^{M}a_{i}\frac{\partial f_{i}^{*}}{\partial\boldsymbol{y}_{k}}+\sum_{j=1}^{N}b_{j}\frac{\partial g_{j}^{*}}{\partial\boldsymbol{y}_{k}}\right),$
where $a_{i}=\sum_{j=1}^{N}C_{ij}\Pi^{*}_{ij}$, and $b_{j}=\sum_{i=1}^{M}C_{ij}\Pi^{*}_{ij}$. The main computation burden is the second term. To compute $\left(\sum_{i}a_{i}\frac{\partial f_{i}^{*}}{\partial\boldsymbol{y}_{k}}+\sum_{j}b_{j}\frac{\partial g_{j}^{*}}{\partial\boldsymbol{y}_{k}}\right)$, a common practice is to solve the linear system (34) to get all $\frac{\partial f^{*}}{\partial\boldsymbol{y}_{k}}$ and $\frac{\partial g^{*}}{\partial\boldsymbol{y}_{k}}$, which is computationally expensive because the right-hand-side of (34) is a third-order tensor $\mathcal{R}$.
Although we do not have (29) as $\boldsymbol{a}$ and $\boldsymbol{b}$ are no longer the marginal constraints of $\Pi^{*}$, we can yet apply the same technique to reduce the computation costs by solving the following matrix-vector form of linear system only once $H(\Pi^{*})\boldsymbol{r}=\left[\begin{array}[]{ccccccccc}\boldsymbol{a}\\ \boldsymbol{b}\end{array}\right]\in\mathbb{R}^{(M+N)}$.
That is, we only need to solve for one time the matrix-vector form of the linear system, for the column vector $\boldsymbol{r}$. This can be efficiently solved by conjugate gradient method with early stopping.
Notice that $\left[\begin{array}[]{ccccccccc}\boldsymbol{a}^{\top}&\boldsymbol{b}^{\top}\end{array}\right]\left[\begin{array}[]{ccccccccc}\mathbbm{1}_{M}\\ -\mathbbm{1}_{N}\end{array}\right]=\mathbb{0}_{M+N}$, and $\left[\begin{array}[]{ccccccccc}\mathbbm{1}_{M}\\ -\mathbbm{1}_{N}\end{array}\right]$ is in the kernel space of $H(\Pi^{*})$ as proved in Lemma 6, so $\left[\begin{array}[]{ccccccccc}\boldsymbol{a}\\ \boldsymbol{b}\end{array}\right]$ is in the span of $H(\Pi^{*})$, therefore the linear system above always has a solution.
@@
$\sum_{i}a_{i}\frac{\partial f_{i}^{*}}{\partial\boldsymbol{Y}}+\sum_{j}b_{j}\frac{\partial g_{j}^{*}}{\partial\boldsymbol{Y}}=\left[\begin{array}[]{ccccccccc}\boldsymbol{a}^{\top}&\boldsymbol{b}^{\top}\end{array}\right]\left[\begin{array}[]{ccccccccc}\frac{d\boldsymbol{f}^{*}}{d\boldsymbol{Y}}\\ \frac{d\boldsymbol{g}^{*}}{d\boldsymbol{Y}}\end{array}\right]$ $=\boldsymbol{r}^{\top}H(\Pi^{*})\left[\begin{array}[]{ccccccccc}\frac{d\boldsymbol{f}^{*}}{d\boldsymbol{Y}}\\ \frac{d\boldsymbol{g}^{*}}{d\boldsymbol{Y}}\end{array}\right]=\boldsymbol{r}^{\top}\mathcal{R}.$

### v2

@@ 3.2 Analytical computation of the gradients
Instead of solving the above linear system to evaluate (14a), we observe from the marginal constraints that
@@
$\left[\begin{array}[]{ccccccccc}\frac{1}{M}\mathbb{1}^{\top}_{M}&\frac{1}{N}\mathbb{1}^{\top}_{N}\end{array}\right]\mathbf{H}^{*}=\left(\frac{1}{M}+\frac{1}{N}\right)\left[\begin{array}[]{ccccccccc}\bm{\mu}^{\top}&\bm{\nu}^{\top}\end{array}\right].$ (27)
@@
Then, multiplying both sides of (24) by $\left[\begin{array}[]{ccccccccc}\frac{1}{M}\mathbb{1}^{\top}_{M}&\frac{1}{N}\mathbb{1}^{\top}_{N}\end{array}\right]$, we obtain
@@
$\left[\begin{array}[]{ccccccccc}\frac{1}{M}\mathbb{1}^{\top}_{M}&\frac{1}{N}\mathbb{1}^{\top}_{N}\end{array}\right]\mathbf{H}^{*}\left[\begin{array}[]{ccccccccc}\frac{d\mathbf{f}^{*}}{d\mathbf{y}_{k}}\\ \frac{d\mathbf{g}^{*}}{d\mathbf{y}_{k}}\end{array}\right]$ $=\left[\begin{array}[]{ccccccccc}\frac{1}{M}\mathbb{1}^{\top}_{M}&\frac{1}{N}\mathbb{1}^{\top}_{N}\end{array}\right]\left[\begin{array}[]{ccccccccc}\mathbf{e}_{k}(\mathbf{B}_{k}\mathbb{1}_{N})^{\top}\\ \mathbf{B}_{k}^{\top}\end{array}\right].$
Hence, the derivative of $\mathrm{OT}_{\epsilon}(\mathbf{C},\bm{\mu},\bm{\nu})$ with respect to source data $\mathbf{y}_{k}$ is
$\frac{d\mathrm{OT}_{\epsilon}(\mathbf{C},\bm{\mu},\bm{\nu})}{d\mathbf{y}_{k}}=\left[\begin{array}[]{ccccccccc}\bm{\mu}^{\top}&\bm{\nu}^{\top}\end{array}\right]\left[\begin{array}[]{ccccccccc}\frac{d\mathbf{f}^{*}}{d\mathbf{y}_{k}}\\ \frac{d\mathbf{g}^{*}}{d\mathbf{y}_{k}}\end{array}\right]$
$=\frac{1}{\frac{1}{M}+\frac{1}{N}}\left[\begin{array}[]{ccccccccc}\frac{1}{M}\mathbb{1}^{\top}_{M}&\frac{1}{N}\mathbb{1}^{\top}_{N}\end{array}\right]\left[\begin{array}[]{ccccccccc}\mathbf{e}_{k}(\mathbf{B}_{k}\mathbb{1}_{N})^{\top}\\ \mathbf{B}_{k}^{\top}\end{array}\right]=\sum_{j=1}^{N}\frac{\partial C_{kj}}{\partial\mathbf{y}_{k}}\Pi^{*}_{kj}.$
@@
If we define the third-order tensor $\mathcal{B}$ by stacking the matrices $\{\mathbf{B}_{k}\}_{k=1}^{M}$, that is the $k$-th component $\mathcal{B}_{k\cdot\cdot}=\mathbf{B}_{k}$, then the linear system (24) can be further vectorized as
@@
$\mathbf{H}^{*}\left[\begin{array}[]{ccccccccc}\frac{d\mathbf{f}^{*}}{d\mathbf{Y}}\\ \frac{d\mathbf{g}^{*}}{d\mathbf{Y}}\end{array}\right]=\mathcal{R},\ \ \text{with}\quad\mathcal{R}:=\left[\begin{array}[]{ccccccccc}\text{diag}(\mathcal{B}\cdot\mathbb{1}_{N})\\ \mathcal{B}^{\top}\end{array}\right]\in\mathbb{R}^{(M+N)\times M\times d}$ (32)
@@
where $\text{diag}(\mathcal{B}\cdot\mathbb{1}_{N})$ is a third-order tensor, with $\text{diag}(\mathcal{B}\cdot\mathbb{1}_{N})_{kks}=(\mathcal{B}\cdot\mathbb{1}_{N})_{ks}$ and zeros for the other entries, and $\mathcal{B}^{\top}$ is the transpose of permutation of indices defined by $\left(\mathcal{B}^{\top}\right)_{ijk}=\mathcal{B}_{kij}$. Hence, by plugging (32) back to (14b), the derivative of $\text{OT}_{\epsilon}(\mathbf{C},\bm{\mu},\bm{\nu})$ with respect to the source data $\mathbf{Y}$ in the tensor form is equal to $\frac{d\mathrm{OT}_{\epsilon}(\mathbf{C},\bm{\mu},\bm{\nu})}{d\mathbf{Y}}=\mathcal{B}\cdot\mathbb{1}_{N}$. ∎
@@
This theorem not only implies that automatic differentiation is unnecessary for computing the gradient of EOT distance but also that solving the large linear system (32) for $\frac{d\mathbf{f}^{*}}{d\mathbf{y}_{k}}$ and $\frac{d\mathbf{g}^{*}}{d\mathbf{y}_{k}}$ is not needed. The analytical expression of the gradient follows directly from the cost matrix $\mathbf{C}_{\theta}$ and optimal coupling matrix $\bm{\Pi}^{*}$ from the Sinkhorn algorithm, which can significantly speed up the computation. This theorem provides a different approach to derive the same analytical result other than the Danskin’s theorem. We emphasize that the gradient formula (16) is generic and applies to any choice of cost matrix $\mathbf{C}$ without requiring $\bm{\mu}$ and $\bm{\nu}$ to be uniform. In particular, if the Sinkhorn iteration stops early and the coupling matrix $\widehat{\bm{\Pi}}$ is suboptimal, the formula is still exact for $\frac{d\mathrm{OT}_{\epsilon}(\mathbf{C},\tilde{\bm{\mu}},\tilde{\bm{\nu}})}{d\mathbf{Y}}$ given suboptimal coupling matrix $\widehat{\bm{\Pi}}$ and associated marginals $\tilde{\bm{\mu}}$ and $\tilde{\bm{\nu}}$, where $\sum_{j}\widehat{\Pi}_{ij}=\widetilde{\mu}_{i}$ and $\sum_{i}\widehat{\Pi}_{ij}=\widetilde{\nu}_{j}$.
@@
Gradient of $\widetilde{\text{OT}}_{\epsilon}(\mathbf{C},\bm{\mu},\bm{\nu})$. Unlike Danskin’s approach, the above techniques also applies for computing the gradient of the Sinkhorn distance $\widetilde{\text{OT}}_{\epsilon}(\mathbf{C},\bm{\mu},\bm{\nu})$. Although it does not yield an analytical expression, it helps significantly reduce the computational cost by avoiding directly solving the linear system (32), which is costly since the right-hand-side is a large third-order tensor $\mathcal{R}$. Specifically, we compute the derivative with respect to the source data $\mathbf{Y}$, whereas prior works, such as [22], have focused on computing the derivative with respect to the measures $\bm{\mu}$ or $\bm{\nu}$. This distinction represents the following novel result.
@@
The gradient of Sinkhorn distance in (11) with respect to the source data $\mathbf{Y}$ is
$\frac{d\widetilde{\mathrm{OT}}_{\epsilon}(\mathbf{C},\bm{\mu},\bm{\nu})}{d\mathbf{Y}}=\left(\mathcal{B}-\frac{\mathcal{A}}{\epsilon}\right)\cdot\mathbb{1}_{N}+\frac{1}{\epsilon}\bm{r}^{\top}\mathcal{R},$ (33)
where $\mathcal{B}$ and $\mathcal{R}$ are defined in (32), the third-order tensor $\mathcal{A}$ is $\mathcal{A}_{ksj}=\frac{\partial C_{kj}}{\partial(\mathbf{y}_{k})_{s}}C_{kj}\Pi^{*}_{kj}$. The vector $\bm{r}$ is the solution of the linear system, $\mathbf{H}^{*}\bm{r}=\left[\begin{array}[]{ccccccccc}\bm{a}\\ \bm{b}\end{array}\right]$, where vectors $\bm{a}$ and $\bm{b}$ have entries $a_{i}=\sum_{j=1}^{M}C_{ij}\Pi^{*}_{ij}$ and $b_{j}=\sum_{i=1}^{N}C_{ij}\Pi^{*}_{ij}$.
@@
From (11) and (17), the derivative of Sinkhorn distance with respect to the data $\mathbf{y}_{k}$ is
$\frac{d\widetilde{\mathrm{OT}}_{\epsilon}(\mathbf{C},\bm{\mu},\bm{\nu})}{d\mathbf{y}_{k}}$ $=\sum_{j=1}^{N}\left(1-\frac{C_{kj}}{\epsilon}\right)\Pi^{*}_{kj}\frac{\partial C_{kj}}{\partial\mathbf{y}_{k}}+\frac{1}{\epsilon}\left(\sum_{i=1}^{M}a_{i}\frac{\partial f_{i}^{*}}{\partial\mathbf{y}_{k}}+\sum_{j=1}^{N}b_{j}\frac{\partial g_{j}^{*}}{\partial\mathbf{y}_{k}}\right),$
where $a_{i}=\sum_{j=1}^{N}C_{ij}\Pi^{*}_{ij}$, and $b_{j}=\sum_{i=1}^{M}C_{ij}\Pi^{*}_{ij}$. The main computation burden is the second term. To compute $\left(\sum_{i}a_{i}\frac{\partial f_{i}^{*}}{\partial\mathbf{y}_{k}}+\sum_{j}b_{j}\frac{\partial g_{j}^{*}}{\partial\mathbf{y}_{k}}\right)$, a common practice is to solve the linear system (32) to get all $\frac{\partial f^{*}}{\partial\mathbf{y}_{k}}$ and $\frac{\partial g^{*}}{\partial\mathbf{y}_{k}}$, which is computationally expensive because the right-hand-side of (32) is a third-order tensor $\mathcal{R}$.
Although we do not have (27) as $\bm{a}$ and $\bm{b}$ are no longer the marginal constraints of $\bm{\Pi}^{*}$, we can yet apply the same technique to reduce the computation costs by solving the following matrix-vector form of linear system only once $\mathbf{H}^{*}\bm{r}=\left[\begin{array}[]{ccccccccc}\bm{a}\\ \bm{b}\end{array}\right]\in\mathbb{R}^{(M+N)}$. That is, we only need to solve for one time the matrix-vector form of the linear system, for the column vector $\bm{r}$. This can be efficiently solved by conjugate gradient method with early stopping.
Notice that $\left[\begin{array}[]{ccccccccc}\mathbb{1}^{\top}_{M}&-\mathbb{1}^{\top}_{N}\end{array}\right]\left[\begin{array}[]{ccccccccc}\bm{a}\\ \bm{b}\end{array}\right]=0$, and $\text{span}\left(\left[\begin{array}[]{ccccccccc}\mathbb{1}_{M}\\ -\mathbb{1}_{N}\end{array}\right]\right)$ is the kernel of $\mathbf{H}^{*}$ as proved in Theorem 6, so $\left[\begin{array}[]{ccccccccc}\bm{a}\\ \bm{b}\end{array}\right]$ is in the span of $\mathbf{H}^{*}$, therefore the linear system above always has a solution.
@@
$\sum_{i}a_{i}\frac{\partial f_{i}^{*}}{\partial\mathbf{Y}}+\sum_{j}b_{j}\frac{\partial g_{j}^{*}}{\partial\mathbf{Y}}=\left[\begin{array}[]{ccccccccc}\bm{a}^{\top}&\bm{b}^{\top}\end{array}\right]\left[\begin{array}[]{ccccccccc}\frac{d\mathbf{f}^{*}}{d\mathbf{Y}}\\ \frac{d\mathbf{g}^{*}}{d\mathbf{Y}}\end{array}\right]$ $=\bm{r}^{\top}\mathbf{H}^{*}\left[\begin{array}[]{ccccccccc}\frac{d\mathbf{f}^{*}}{d\mathbf{Y}}\\ \frac{d\mathbf{g}^{*}}{d\mathbf{Y}}\end{array}\right]=\bm{r}^{\top}\mathcal{R}.$